If a non-cyclic alkane shows a molecular ion peak at m/z 3.80 × 102, what is the chemical formula?
Answer #1

A non-cyclic alkane shows a molecular ion peak at m/z 3.80 × 102= 380

Let n be the number of carbons in the alkane. Then the number of hydrogens is 2n+2. The combined weight of the carbons plus the combined weight of the hydrogens must add up to 380. So we have:

12xn + 1x(2n+2) = 380
12n+2n+2 = 380

n = 378/14 = 27
we get n = 27 and 2n+2 = 56

Therefore the molecular formula of non-cyclic alkane is C27H56.